OpenStudy (anonymous):

Can anyone help me with solving this differential equaton: y^4-y^3*y''=1 where y=y(x)

OpenStudy (jamesj):

Tricky ... in that it needs some sort of trick of substitution, which I haven't seen yet. In the meantime, here's the exact solution: http://www.wolframalpha.com/input/?i=y%5E4-y%5E3*y%27%27%3D1

OpenStudy (jamesj):

I observe that if we write \( u = (2y)^2 \), then from the Wolfram solution, we have \[ u = (c_1 f - 1/f)^2 - 4f^2 \] where \( f(x) = e^{x + c_2} \) and hence satisfies the equation \( f'/f = 1 \). In my approach, the next step is to try and reverse engineer the method to solve the original equation. If I get any further with this later today I'll let you know.

OpenStudy (anonymous):

Nothing? Because this to me seems like a real tough stuff. I was trying to approach to it by trying to regroup the equation: we can see, that (y^4)'=4*y^3*y' , although it has y^3, it has first derivative y', but we need y'', so this approach got me nowhere.

OpenStudy (anonymous):

And basically saying \[y^3(y-y'')=1\] gives us nothing.

OpenStudy (anonymous):

Finally thought of a way to get some kind of solution: \[y^4-y^3*y''=1\]\[y''=(y^4-1)/y^3\] multiply both sides with 2y' \[2y'*y''=2y'(y-1/y^3)\]\[2y'*y''=(y'^2)'\]\[(y'^2)'=2y'(y-1/y^3)\] Integrate both sides: \[(y')^2=2 \int\limits y'*(y-1/y^3)dy\]\[(y')^2=y^2+1/y^2+C_1\] Now I guess we take square root of both sides and integrate one time. Of course solving the equation like this, we can get solution in form of x=x(y), but for now I cannot think of anything else.

OpenStudy (jamesj):

You've got it. \[ y^2 (y')^2 = 1 + y^4 + C1.y^2 \] and thus looking at the positive root, \[ \frac{y}{\sqrt{1+C_1.y^2 + y^4}} dy = dx \] Now substitute u = y^2, du = 2y dy \[ \frac{du}{2\sqrt{(u + C_1/2)^2 + (1-C_1^2/4)}} = dx \] which you can now integrate \[ \ln(\sqrt{ (u + C_1/2)^2 + (1-C_1^2/4) } + (u + C_1/2)) = x + C_2 \] and this you can now unwind to solve for u, then y.

OpenStudy (jamesj):

...I dropped the 1/2

OpenStudy (anonymous):

Hmm.. why only positive root?

OpenStudy (jamesj):

You need both obviously. I just wanted to show it's straightforward.